OpenStudy (anonymous):

I'm failing chem 1 i think i'm having the most problems in finding moles and using stochiometry I really don't get how to get moles in the first place either

OpenStudy (abb0t):

Ok, I am going to start with a simple example to help clarify things for you: Example question: How many moles of oxygen, O, is required to burn 2.40 moles of ethane (C2H6) grams.

OpenStudy (abb0t):

First of all, you should know that anytime that you react something with oxygen and heat, you get carbon dioxide and water as products. I don't think you need to, but for practice (since you say you're "failing") it might be wise to start with a balanced equation. The equation for this is:

OpenStudy (abb0t):

\[2 C_2H_6+7O_2 \rightarrow 4 CO_2 + 6 H_2O\] Now, your first step is to make a PLAN out the problem. Identify the GIVEN and the WANTED quantities and PER/PATH

OpenStudy (abb0t):

GIVEN: 2.40 mol ethane (C2H6) WANTED Mol O2 \[PER = \frac{ 7 mol. O_2 }{ 2 mol. C_2H_6 }\] Therefore, the pathway you are taking is: \[mol. C_2H_6 \rightarrow mol. O_2\]

OpenStudy (abb0t):

The conversion factor comes from the PER relationship between the given and the wanted quantities, using their coefficients (7 & 2).

OpenStudy (abb0t):

So, using dimensional analysis to solve for what is being asked for, you have: \[2.40 mol. C_2H_6 \times \frac{ 7 mol. O_2 }{ 2 mol. C_2H_6 } = 8.40 mol. O_2\]

OpenStudy (abb0t):

Notice how I got the answer, it's very similar to fractions! The mol of ethane both cancel out! It's the same as solving for "y" in this example: \[2.40x \times \frac{ 7y }{ 2x }=8.40y\] The two x's cancel out and you're left with:\[2.40 \times \frac{ 7y }{ 2 }= 8.40y\]

OpenStudy (anonymous):

thats actually making a lot more sense then the teacher ever did

OpenStudy (abb0t):

where y = moles of O2

OpenStudy (abb0t):

For exams and quizzes, it will probably be expected that you mastered stoichiometry so you are expected to know how to balance a chemical equation. Here's ONE more example that I will let you solve alone. Same difficulty, but much simpler molecule: Question: Ammonia is simply two elements. How many moles of HYDROGEN are needed to produce 2.40 moles of ammonia??? It might help to write down what I told you to before, start by identifying: 1. GIVEN 2. WANTED and plan out the PER/ PATH.

OpenStudy (abb0t):

The PER/PATH might not be needed for this problem, BUT you will use it for further problems such as trying to find mass from moles, or grams of X to grams of Y and so forth. Best of luck!

OpenStudy (anonymous):

the only part i don't get is where you got the 8.40mol.02 from

OpenStudy (abb0t):

That's the amount of moles. It's like I explained. It's like algebra. You know that \[\frac{ 4 }{ 2 } = 2\] what it means is basically dividing. What I did was: \[2.40 \div 2 \times 7 = 8.40\]

OpenStudy (anonymous):

and can you give the written equation for the problem you gave me not balanced i just need the equation

OpenStudy (abb0t):

I don't want to give you the equation because during an exam you may or may not be given the equation. I believe that your first week(s) of chemistry, you are taught to memorize the elements, their names, and also memorize some polyatomic ions and names of compounds and such... I will give you a hint, look @ the question, it says "Ammonia is simply two elements". The formula for ammonia is: NH3. Therefore, to make ammonia (NH3), the two elements must be: _____________ & _______________

OpenStudy (anonymous):

NH & H2 ?

OpenStudy (abb0t):

Not quite. nitrogen exists naturally as a diatom, so it would NOT be NH. However, you ARE correct with hydrogen.

OpenStudy (anonymous):

so would it be N2H & H2 ?

OpenStudy (abb0t):

sorry, i mean't diatom under standard conditions*

OpenStudy (abb0t):

And no, N2H would I think be a compound, not an element. When they say element, it means the molecule by itself w/o anything attached to it. Hydrogen exists as a diatom, too. That's why it's H2, and nitrogen similarly, is N2

OpenStudy (anonymous):

would the equation be 2NH3--->N2+ 3H2

OpenStudy (abb0t):

Correct, but the problem said "forming". And what you have there is "breaking" so instead, it might be best to re-write it like this: \[N_2 + 3H_2 \rightarrow 2NH_3\]

OpenStudy (anonymous):

yeah i meant reversed but that is the right way?

OpenStudy (abb0t):

Yes, just re-write it how I did. I know it sounds like it doesn't matter and its too picky, but chemistry is universal and people around the world must be able to understand. And if you write it like that someone else might think of it as decomposition of ammonia even though you meant opposite. Just want to clarify this. I am not sure how picky your professor is, but its best to avoid any confusion to avoid loss of easy points.

OpenStudy (anonymous):

PER= 1mil. N2/ 3 mol. H2

OpenStudy (anonymous):

*mol

OpenStudy (abb0t):

Not quite, you are being asked how many HYDROGEN. Identify what you are given first, it might clarify things.

OpenStudy (abb0t):

The problem asks nothing about nitrogen. It's only referring to hydrogen and ammonia.

OpenStudy (anonymous):

i think i had it upside down I meant 3mol.H2/ 1 Mol NH3

OpenStudy (abb0t):

You almost have it, but if you look back at your balanced euqation, how many moles of ammonia are there? is there really only one?

OpenStudy (anonymous):

3molH2/ 2mol NH3

OpenStudy (abb0t):

Correct!

OpenStudy (anonymous):

Yay :D

OpenStudy (abb0t):

You're almost done! Now use the given info to figure out the moles of hydrogen

OpenStudy (anonymous):

so 2.40 mol NH3x3H2/2NH3= 3.60 mol. H2

OpenStudy (abb0t):

Good job. You're correct.

OpenStudy (anonymous):

*3molH2/2molNH3

OpenStudy (anonymous):

:D thank you so much for helping me and walking me through this I actually understand it for the most part now I have to go but thatnk you so much!